# SLO-aware GPU Frequency Scaling for Energy Efficient LLM Inference Serving

### v1

@@ IV-A Generation Length Prediction
IV-B KV usage & Batch Size projection
@@
To keep track of scheduled requests, a Scoreboard is kept, containing the following metadata for each query ($q_{i}$) in the batch: i) the iteration at which the query was scheduled ($s_{i}$), ii) the input length of the query ($|q_{i}|$) and iii) its predicted generation length ($\widehat{|r_{i}|}$). Based on these values, we can determine that a query will complete its generation process at iteration $s_{i}+\widehat{|r_{i}|}$, i.e., $\widehat{|r_{i}|}$ iterations after it was initially scheduled. By repeating this process for each entry in the Scoreboard, we can identify the final iteration of all scheduled queries, allowing us to generate a vector $B=[B_{1},B_{2},...,B_{j},...,B_{n}]$, where $B_{j}$ is the projected batch size at each future iteration $j\in[1,n]$ and $n$ is the iteration at which the last query will be completed.
@@
Last, when a query terminates , its entry is struck from the Scoreboard, signalling the deallocation of all associated KV cache blocks. The introduction of this component allows us to project the usage of the KV cache and the batch size into the future, under the assumption that no new requests will arrive. We measure an average latency of $<2$ ms for this component.
Projections for new queries: When a new query $q_{a}$ is submitted and before it is scheduled, its metadata are “virtually” appended to the Scoreboard with $s_{a}$ being set to the current iteration $k$, thus, allowing us to assess how future KV cache usage and batch size would be affected by the immediate scheduling of the request. If the query is indeed scheduled by throttLL’eM’s scheduler (Sec.IV-C), this “virtual” change is committed to the Scoreboard, otherwise, it is rolled back.
IV-C LLM Query Scheduler
@@
IV-C1 Performance Prediction Model
The Performance Prediction Model ($\mathcal{M}$) predicts the throughput of the system in terms of iterations per second (IPS). Building on the high correlation of throughput to KV cache usage, GPU frequency, and batch size demonstrated in Sec. III-A1 and Sec. III-B, the model leverages these metrics as input features to accurately forecast system performance at the iteration level.
Training data collection: Capturing the entire design space for training the predictive model is prohibitive due to i) the fine granularity of GPU frequencies (15MHz steps), ii) the extensive range of possible values for KV block usage, and iii) the large batch sizes, which can reach up to 512 instances in few-parameter models with a high degree of parallelism, or on GPU infrastructures with higher memory capacity [4]. To address this challenge, we employ a systematic sampling approach to acquire a representative subset of the design space. Specifically, we employ a request generator that profiles the LLM inference engine across different batch sizes by spawning queries with predefined generation lengths, ensuring maximum utilization of the KV cache at the final iteration. For each batch size, the generator spawns a number of requests equal to the batch size to cover the entire KV cache usage range, ensuring that the edges of the profiling space are present in the dataset. A monitoring agent is deployed to continuously log KV cache usage, GPU frequency, batch size, and the experienced IPS every second. During this process, the GPU frequency is randomly changed (for high frequency range coverage) after each measurement is completed and before starting a new one, in order to ensure that the frequency remains consistent for the duration of each measurement. This data collection process is repeated for all supported tensor parallelism (TP) levels, generating a dataset for each LLM that captures the relationship between the input variables and IPS.
Model selection & training: throttLL’eM adopts a machine learning (ML) approach to model the relationship between system metrics and IPS. Given that all the input features are upper and lower bounded for a given LLM, we employ a Gradient Boosted Decision Tree [20, 10] as our ML model, which has been shown to provide accurate interpolation predictions for bounded datasets with multiple interacting features [33]. The model is lightweight ($\approx 3$ ms inference latency on CPU), meeting the design constraint of low inference latency. This is vital not only for enabling iteration-level predictions but also because the model resides in the critical path of the Scheduler. We use the following features as input to the model: i) engine size, ii) batch size, iii) KV cache usage and iv) GPU frequency, training it to predict the throughput (IPS).
IV-C2 Query Scheduling & Admission Control
@@
If all checks are successful, the newly arrived request is scheduled and any changes to the Scoreboard are committed. Otherwise, the request is queued and changes are rolled back. If a new request cannot be scheduled without violation of its own E2E SLO, but does not cause violations, it is scheduled but marked as “lost”, signalling the Scheduler to ignore it during future SLO validations.
IV-D LLM autoscaling mechanism
Effective autoscaling is crucial for optimizing both performance and energy efficiency in LLM inference systems. As shown in Sec. III-C, tensor parallelism (TP) forms the most efficient approach for scaling the LLM engine to multiple GPUs on the same node. To this end, throttLL’eM’s autoscaling mechanism relies on managing the TP of the LLM engine to match the incoming request rate (RPS), ensuring that the system operates efficiently under varying loads.
Autoscaling is driven by a 10-second interval monitoring agent that determines the minimum engine size required to handle the current workload. This decision is based on pre-characterized performance profiles, which define the maximum load each engine can sustain without introducing long tail latencies (cf. Sec. V-A). To mitigate the significant provisioning latency ($>$20s) of new inference servers, we employ “shadow instancing[12, 18], thus, masking switching overheads. This involves a “warm-up” phase, where new requests are served by the current engine until the new one becomes operational. Afterwards, a “transition” phase occurs, during which the old engine gracefully shuts down, while the new engine takes over, serving all new requests. During shadow instancing, both engines run simultaneously, consuming energy without providing peak performance capacity until fully loaded.
@@ IV-D LLM autoscaling mechanism
IV-E GPU Frequency Throttling Controller
The throttling controller is triggered upon admission of a new query to the inference engine. Its goal is to identify the minimum GPU frequency that adheres to the target SLOs, unlike the Scheduler, which validates SLO compliance at maximum frequency. Since this component is invoked subsequent to successful SLO validation at maximum frequency, the existence of at least one SLO-compliant GPU frequency is guaranteed. The controller receives the $KV$ and $B$ vectors from the KV and Batch projection mechanism, performing a binary search over the GPU frequency range. This involves: 1) Recalculating $T$, $T^{\prime}$, and $\widehat{T_{R}}$ for each frequency; and 2) Checking TBT and E2E SLO compliance using the same procedure as the Scheduler. This process determines the minimum SLO-satisfying frequency. If a “lost” request is present in the current batch, the binary search is bypassed, and the maximum GPU frequency is selected in an attempt to meet the “lost” SLO. We measure the combined latency of the Scheduler and GPU frequency throttling components at 35 ms under heavy load.
IV-F Discussion
@@
throttLL’eM and TTFT optimization: throttLL’eM does not explicitly optimize the prefill phase, a compute-intensive part of LLM inference. While the prefill phase typically constitutes a smaller proportion of overall latency compared to the generation phase [43], it still represents a potential performance bottleneck. Increasing GPU frequency for a single iteration (upon scheduling a new query) would be an inefficient way to manage TTFT due to the frequency switching overhead (avg. 200 ms) compared to the short prefill phase duration (avg. $\approx$175 ms). The Splitwise framework [43] offers a promising alternative by decoupling the prefill and generation stages onto separate GPUs, allowing for independent optimization.
V Evaluation
V-A Experimental Setup
HW/SW stack: We deploy throttLL’eM on a p4d.24xlarge AWS EC2 instance equipped with 8×8$\times$8 × NVIDIA A100 40GB GPUs [5]. For the software stack, we use NVIDIA’s Triton server [40] and TensorRT-LLM backend [39] which provides high-performance inference on NVIDIA GPUs. We enable typical LLM optimizations, i.e., i) inflight fused batching [65], ii) paged KV caching[28] and iii) Flash Attention[16].
TABLE II: Performance profiles of examined LLMs.
5
@@ V-A Experimental Setup
Load generation: Workload is generated using invocation traces from Azure [43], analyzed in Sec. III-D. Since the trace has a peak RPS of $\approx 8.25$, we right-scale the invocation rate to match the maximum load of the respective evaluated engine (reported in Table II). As the trace does not provide the actual context of input and output (due to GDPR regulations), we generate synthetic queries that match the prompt and generation lengths of each item in the dataset.
V-B Performance Prediction Model Evaluation
@@
Table III reports the $R^{2}$ score, Mean Absolute Error (MAE) and Mean Absolute Percentage Error (MAPE) of throttLL’eM performance prediction model ($\mathcal{M}$) (Sec. IV-C1), for each one of the examined engines and under two different train/test splits. Using a $90/10$ split, $\mathcal{M}$ accurately predicts target IPS with $R^{2}$ scores $\geq 0.97$, often exceeding $0.98$. MAPE is kept below $5.8$% for all configurations, reaching a low of $2.8$% for Llama2-13B when running on a single GPU. MAE, providing a more tangible representation of the results, indicates mispredictions under $1$ IPS on average for all configurations. With a $10/90$ train/test split, prediction performance remains robust, featuring $R^{2}$ scores above $0.96$, MAE below $1.01$ IPS, and a MAPE increase of $0.7$% at most. These results demonstrate that high accuracy is achievable even with sparse training sets.
V-C KV cache and Batch size Projection Mechanism Evaluation
@@
We evaluate the KV cache usage and batch size projection component (Sec. IV-B) using custom generated micro-traces. Each micro trace sets a different GPU frequency and includes queries with random prompt and generation lengths. We spawn all queries simultaneously which trigger throttLL’eM’s projection and prediction mechanisms, generating vectors $KV$, $B$ and $\widehat{T_{R}}$ for the specified frequency, which contains the estimated remaining time to reach all future iterations. At the same time, the performance monitoring endpoint of Triton is continuously polled, logging the actual batch size, KV cache usage and arrival time at each iteration. Figures 7a7b show the error distributions between actual and predicted batch size and KV cache usage respectively, while Fig. 7c plots the accumulated drift, i.e., how much predicted times (elements of $\widehat{T_{R}}$) diverge from actual arrival times, divided by the number of elapsed iterations. Overall, batch size projection errors are kept to a minimum, averaging an error of 0.19%. The KV projection error is slightly higher while still remaining low at 2.26% on average. Finally, each LLM iteration introduces a relatively small drift (avg. 0.43ms) when compared to the TBT (15 to 30ms) for the same inference engine, shown in Fig. 2c.
@@
TABLE III: Performance Prediction Model Evaluation
7
@@
V-D throttLL’eM end-to-end Evaluation
V-D1 throttLL’eM without autoscaling
@@
Figure 8a shows the distribution of E2E latencies for different engines when Triton[40] and throttLL’eM are deployed on the scaled trace. In almost all cases, with the exception of llama2-13b-TP1, throttLL’eM achieves the specified $p99th$ latency SLO, being on average 1.45 seconds under the deadline. We attribute the inability of throttLL’eM to achieve the E2E SLO for llama2-13b-TP1 to two factors. First, even the default Triton implementation stays just 1.7 seconds below the SLO, hinting that throttling may not be applicable to this situation. Secondly, the very limited number of KV blocks available (120 per Tab. II) means that i) the engine is limited to lower batch size and throughput and ii) requests are often queued to avoid swapping to system memory. The TBT SLO is achieved in all cases, as depicted in Fig. 8b, with the vast majority of TBT values falling well below 200 ms. Outlier TBT values can be attributed to inflight batching overheads, as token generation stalls until the prefill phase of a new request is completed.
throttLL’eM aims to increase the energy efficiency of LLM serving. Fig. 8c examines the distribution of power consumption values encountered by Triton and throttLL’eM under different prediction error levels. Triton displays consistently high power draw for all LLM engines. On the other hand, the proposed solution offers reduced power consumption due to lower GPU frequencies. As throttLL’eM dynamically adjusts frequencies to meet SLOs, a wide range of power values is encountered, including peak power equal to that of Triton when under high system pressure. Overall, throttLL’eM offers 36.3% higher energy efficiency on average with oracle predictions, dropping to 30.0% at 30% prediction error level. The highest increase in energy efficiency was experienced for llama2-13b-TP2, producing up to 44.3% more tokens per Joule than Triton. Overall, energy consumption is reduced by an average of 24.7% and up to 30.7%.
(a) End-to-End latency distribution. The red dotted line denotes the E2E SLO of each model as defined in Table II. Text boxes report the $99^{th}$ E2E latency percentile achieved by the different approaches.
@@
Figure 8: Comparison of E2E, TBT and Power consumption between Triton [40] and throttLL’eM. At each plot, the leftmost box represents Triton, followed by throttLL’eM with error levels of $0\%$, $15\%$, and $30\%$, respectively.
(a) GPU Frequency
(b) Queue time
(c) TTFT
@@
These improvements are enabled by the selection of lower frequencies. The barplots of Figure 9a, show that throttLL’eM selects lower, more efficient frequencies than the default 1410 MHz, ranging between 950 and 1260 MHz on average, depending on the LLM engine and the prediction error of the predictor. Higher prediction errors lead to the selection of higher frequencies, as throttLL’eM increases the predicted generation length, leading to frequency overprovisioning.
Finally, we analyse the overheads imposed by the introduction of our framework. throttLL’eM maintains a queue where queries are held until the SLOs are validated. This introduces additional latency due to the time needed to perform the validation and the possible re-queuing of a request if it is predicted to cause violations. Figure 9b contains representative examples for both situations. Llama3-8b-TP1 exhibits relatively high queuing delays as high request arrival rates stress the Scheduler, which can only schedule one request at a time. Llama2-13b-TP1 also exhibits delays, as the limited KV capacity leads to queuing until sufficient blocks become available. From the user’s perspective queuing delays add to the latency of the prefill phase (TTFT). Figure 9c shows that throttLL’eM increases TTFT when compared to Triton. A significant part of this increase is attributed to queuing, while the rest is due to the lower frequencies selected, which slow down the compute-bound prefill phase [4].
V-D2 throttLL’eM with autoscaling
@@
V-E throttLL’eM Runtime Analysis
@@
In general, throttLL’eM selects lower frequencies when the active engine is under light load. When however encountering high pressure, e.g., during engine up-scaling, the throttling controller applies higher GPU frequencies to achieve the desired performance. In addition, the system still applies the frequencies required to meet the SLOs of individual queries.
@@
Figure 11 features some points of interest. The red stars mark transient E2E SLO violations due to a load increase that cannot be absorbed by the TP2 engine while TP4 comes online. Yellow stars highlight how frequency is reduced when the instance is operating at lower load but cannot be downsized. Finally, green stars mark consecutive switches, as the encountered load remains close to the TP1/TP2 boundary.

### v2

@@ 4.1 Generation Length Prediction
4.2 KV usage & Batch Size projection
@@
To keep track of scheduled requests, a Scoreboard is kept, containing the following metadata for each query ($q_{i}$) in the batch: i) the iteration at which the query was scheduled ($s_{i}$), ii) the input length of the query ($|q_{i}|$) and iii) its predicted generation length ($\widehat{|r_{i}|}$). Based on these values, we can determine that a query will complete its generation process at iteration $s_{i}+\widehat{|r_{i}|}$, i.e., $\widehat{|r_{i}|}$ iterations after it initially scheduled. By repeating this process for each entry in the Scoreboard, we can identify the final iteration of all scheduled queries, allowing us to generate a vector $B=[B_{1},B_{2},...,B_{j},...,B_{n}]$, where $B_{j}$ is the projected batch size at each future iteration $j\in[1,n]$ and $n$ is the iteration at which the last query will be completed.
@@
Last, when a query terminates, its entry is struck from the Scoreboard, signalling the deallocation of all associated KV cache blocks. The introduction of this component allows us to project the usage of the KV cache and the batch size into the future, under the assumption that no new requests will arrive. We measure an average latency of <2<2 ms for this component.
Projections for new queries: When a new query $q_{a}$ is submitted and before it is scheduled, its metadata are “virtually" appended to the Scoreboard with $s_{a}$ being set to the current iteration $k$, thus, allowing us to assess how future KV cache usage and batch size would be affected by the immediate scheduling of the request. If the query is indeed scheduled by throttLL’eM’s scheduler (Sec.4.3), this “virtual" change is committed to the Scoreboard, otherwise, it is rolled back.
4.3 LLM Query Scheduler
@@
4.3.1 Performance Prediction Model
The Performance Prediction Model ($\mathcal{M}$) predicts the throughput of the system in terms of iterations per second (IPS). Building on the high correlation of throughput to KV cache usage, GPU frequency, and batch size, as demonstrated in Sec. 3.1.1 and Sec. 3.2, the model leverages these metrics as input features to accurately forecast system performance at the iteration level.
Training data collection: Capturing the entire design space for training the predictive model is prohibitive due to i) the fine granularity of GPU frequencies (15 MHz steps); ii) the extensive range of possible values for KV block usage; and iii) the large batch sizes, which can reach up to 512 instances in few-parameter models with a high degree of parallelism, or on GPU infrastructures with higher memory capacity [agrawal2023sarathi]. To address this challenge, we employ a systematic sampling approach to acquire a representative subset of the design space. Specifically, we employ a request generator that profiles the LLM inference engine across different batch sizes by spawning queries with predefined generation lengths, ensuring maximum utilization of the KV cache at the final iteration. For each batch size, the generator spawns a number of requests equal to the batch size to cover the entire KV cache usage range, ensuring that the edges of the profiling space are present in the dataset. A monitoring agent is deployed to continuously log KV cache usage, GPU frequency, batch size, and the experienced IPS every second. During this process, the GPU frequency is randomly changed (for high frequency range coverage) after each measurement is completed and before starting a new one, in order to ensure that the frequency remains consistent for the duration of each measurement. This data collection process is repeated for all supported tensor parallelism (TP) levels, generating a dataset for each LLM that captures the relationship between the input variables and IPS.
Model selection & training: throttLL’eM adopts a machine learning (ML) approach to model the relationship between system metrics and IPS. Given that all input features are upper and lower bounded for a given LLM, we employ a Gradient Boosted Decision Tree [friedman2001greedy, xgboost] as our ML model, which has been shown to provide accurate interpolation predictions for bounded datasets with multiple interacting features [malistov2019gradient]. The model is lightweight ($\approx 3$ ms inference latency on CPU), meeting the design constraint of low inference latency. This is vital not only for enabling iteration-level predictions but also because the model resides in the critical path of the Scheduler. We use the following features as input to the model: i) engine size; ii) batch size; iii) KV cache usage; and iv) GPU frequency, training it to predict the throughput (IPS).
4.3.2 Query Scheduling & Admission Control
@@
For queued requests, this admission control process is repeated until the Scheduler decides that they can be safely admitted to the batch without causing violations. If all checks are successful, the newly arrived request is scheduled and any changes to the Scoreboard are committed. Otherwise, the request is queued and changes are rolled back. If a new request cannot be scheduled without violation of its own E2E SLO, but does not cause violations, it is scheduled but marked as “lost", signaling the Scheduler to ignore it during future SLO validations.
4.4 LLM autoscaling mechanism
Effective autoscaling is crucial for optimizing both performance and energy efficiency in LLM inference systems. As shown in Sec. 3.3, tensor parallelism (TP) forms the most efficient approach for scaling the LLM engine to multiple GPUs on the same node. To this end, throttLL’eM’s autoscaling mechanism relies on managing the TP of the LLM engine to match the incoming request rate (RPS), ensuring that the system operates efficiently under varying loads.
Autoscaling is driven by a 10-second interval monitoring agent that determines the minimum engine size required to handle the current workload. This decision is based on pre-characterized performance profiles, which define the maximum load each engine can sustain without introducing long tail latencies (cf. Sec. 5.1). To mitigate the significant provisioning latency (>>20s) of new inference servers, we employ “shadow instancing"[shadowMarcusChow, dhakal2020gslice], thus masking switching overheads. This involves a “warm-up” phase, where new requests are served by the current engine until the new one becomes operational. Afterwards, a “transition” phase occurs, during which the old engine gracefully shuts down, while the new engine takes over, serving all new requests. During shadow instancing, both engines run simultaneously, consuming energy without providing peak performance capacity until fully loaded.
@@ 4.4 LLM autoscaling mechanism
4.5 GPU Frequency Throttling Controller
The throttling controller is triggered upon admission of a new query to the inference engine. Its goal is to identify the minimum GPU frequency that adheres to the target SLOs, unlike the Scheduler, which validates SLO compliance at maximum frequency. Since this component is invoked subsequent to successful SLO validation at maximum frequency, the existence of at least one SLO-compliant GPU frequency is guaranteed. The controller receives the $KV$ and $B$ vectors from the KV and Batch projection mechanism, performing a binary search over the GPU frequency range. This involves: 1) Recalculating $T$, $T^{\prime}$, and $\widehat{T_{R}}$ for each frequency; and 2) Checking TBT and E2E SLO compliance using the same procedure as the Scheduler. This process determines the minimum SLO-satisfying frequency. The selected frequency is applied to all the GPUs serving the LLM instance, thus, minimizing operational complexity and mitigating barrier synchronization issues due to stragglers [yang2020mitigating]. If a “lost” request is present in the current batch, the binary search is bypassed, and the maximum GPU frequency is selected in a best-effort attempt to meet the “lost” SLO. We measure the combined latency of the Scheduler and GPU frequency throttling components at 35 ms under heavy load.
4.6 Discussion
@@
throttLL’eM and TTFT optimization: throttLL’eM does not explicitly optimize the prefill phase, a compute-intensive part of LLM inference. While the prefill phase typically constitutes a smaller proportion of overall latency compared to the generation phase [patel2024splitwise], it still represents a potential performance bottleneck. Increasing GPU frequency for a single iteration (upon scheduling a new query) would be an inefficient way to manage TTFT due to the frequency switching overhead (avg. 200 ms) compared to the short prefill phase duration (avg. $\approx$175 ms). The Splitwise framework [patel2024splitwise] offers a promising alternative by decoupling the prefill and generation stages onto separate instances, allowing for independent optimization.
throttLL’eM scalability: The proposed solution uses TP to distribute LLMs across GPUs, an approach shared with other concurrent works [patel2024splitwise, patel2023polca]. TP offers low latency but requires high performance interconnects and may not scale beyond a single node. As discussed in [agrawal2023sarathi], for larger models which need to be distributed across nodes, PP is often used for inter-node communication, while TP is employed intra-node to take advantage of its increased performance. In this case, the modelling methodology presented still applies within each node but additional components would be required in order to orchestrate execution across stages. throttLL’eM is orthogonal to DDP, used by [patel2023polca, patel2024splitwise] to deploy multiple LLM instances.
throttLL’eM overheads: The proposed solution employs a generation length prediction model. Previous works [qiu2024efficient] show that such models introduce an average latency of 7.67.6ms, which is negligible when compared to the multiple seconds of the inference process. throttLL’eM also requires profiling to generate the performance prediction model: a process that takes up to a day for LLaMa3-70B. However, this is a one-time cost that can be amortized over the lifetime of the service and across multiple instances. The prediction model requires less than a second of training on a multi-core CPU, while inference takes on average 33ms. Overall, when throttLL’eM is deployed, the Scheduler and frequency throttling components add 3535ms to i) forecast KV cache utilization and batch sizes (22ms), ii) check multiple frequencies and schedule a query. Finally, setting a new frequency requires 88ms per GPU, which however occurs after request scheduling, off the critical path.
5 Evaluation
5.1 Experimental Setup
HW/SW stack: We deploy throttLL’eM on a p4d.24xlarge AWS EC2 instance equipped with $8\times$ NVIDIA A100 40GB GPUs [aws-ec2-p4dspecs]. For the software stack, we use NVIDIA’s Triton server [triton-inference-server] and TensorRT-LLM backend [tensorRT-LLM-backend] which provides high-performance LLM inference on NVIDIA GPUs. We enable typical LLM performance optimizations, such as: i) inflight fused batching [gyeong2022orca]; ii) paged KV caching[kwon2023pagedattention]; and iii) Flash Attention[dao2022flashattention].
Table 2: Performance profiles of examined LLMs.
@@ 5.1 Experimental Setup
Load generation: Workload is generated using invocation traces from Azure [patel2024splitwise], analyzed in Sec. 3.4. Since the trace has a peak RPS of $\approx 8.25$, we right-scale the invocation rate to match the maximum load of the respective evaluated engine (reported in Table 2). As the trace does not provide the actual context of input and output due to data protection regulations, we generate synthetic queries that match the prompt and generation lengths of each item in the dataset.
5.2 Performance Prediction Model Evaluation
@@
Table 3 reports the $R^{2}$ score, Mean Absolute Error (MAE) and Mean Absolute Percentage Error (MAPE) of the performance prediction model ($\mathcal{M}$) employed by throttLL’eM (Sec. 4.3.1), for each one of the examined engines and under two different train/test splits. Using a $90/10$ split, $\mathcal{M}$ accurately predicts target IPS with $R^{2}$ scores $\geq 0.97$, often exceeding $0.98$. MAPE is kept below 5.85.8% for all configurations, reaching a low of 2.82.8% for Llama2-13B when running on a single GPU. MAE, providing a more tangible representation of the results, indicates mispredictions under $1$ IPS on average for all configurations. With a $10/90$ train/test split, prediction performance remains robust, featuring $R^{2}$ scores above $0.96$, MAE below $1.01$ IPS, and a MAPE increase of 0.70.7% at most. These results demonstrate that high accuracy is achievable even with sparse training sets.
5.3 KV cache and Batch size Projection Mechanism Evaluation
@@
We evaluate the KV cache usage and batch size projection component (Sec. 4.2) using custom generated micro-traces. Each micro-trace sets a different GPU frequency and includes queries with random prompt and generation lengths. We spawn all queries simultaneously which trigger throttLL’eM’s projection and prediction mechanisms, generating vectors $KV$, $B$ and $\widehat{T_{R}}$ for the specified frequency, which contains the estimated remaining time to reach all future iterations. At the same time, the performance monitoring endpoint of Triton is continuously polled, logging the actual batch size, KV cache usage and arrival time at each iteration. Figures 7a7b show the error distributions between actual and predicted batch size and KV cache usage respectively, while Fig. 7c plots the accumulated drift, i.e., how much predicted times (elements of $\widehat{T_{R}}$) diverge from actual arrival times, divided by the number of elapsed iterations. Overall, batch size projection errors are kept to a minimum, averaging an error of 0.19%. The KV projection error is slightly higher while still remaining low at 2.26% on average. These small variations are attributed to the latency experienced between request submission and admission, during which execution progresses, occasionally resulting in off-by-one errors when calculating the final iteration of each request. Finally, each LLM iteration introduces a relatively small average drift of 0.43 ms, compared to the TBT (15 to 30 ms) for the same inference engine, shown in Fig. 2c. This drift is a combined effect of i) the aforementioned errors in KV cache usage and Batch size projections; and ii) estimation errors of the performance prediction model $\mathcal{M}$.
@@
Table 3: Performance Prediction Model Evaluation
@@
5.4 throttLL’eM end-to-end Evaluation
5.4.1 throttLL’eM without autoscaling
@@
Figure 8a shows the distribution of E2E latencies for different engines when Triton[triton-inference-server] and throttLL’eM are deployed on the scaled trace. In almost all cases, with the exception of llama2-13b-TP1, throttLL’eM achieves the specified $p99th$ latency SLO, being on average 1.45 seconds under the deadline. We attribute the inability of throttLL’eM to achieve the E2E SLO for llama2-13b-TP1 to two factors. First, even the default Triton implementation stays just 1.7 seconds below the SLO, hinting that throttling may not be applicable to this situation. Secondly, the very limited number of KV blocks available (120 per Tab. 2) means that i) the engine is limited to lower batch size and throughput and ii) requests are often queued to avoid swapping to system memory. The TBT SLO is achieved in all cases, as depicted in Fig. 8b, with the vast majority of TBT values falling well below 200 ms. Outlier TBT values can be attributed to inflight batching overheads, as token generation stalls until the prefill phase of a new request is completed.
throttLL’eM aims to increase the energy efficiency of LLM serving. Figure 8c examines the distribution of power consumption values encountered by Triton and throttLL’eM under different prediction error levels. Triton displays consistently high power draw for all LLM engines. On the other hand, the proposed solution offers reduced power consumption due to lower GPU frequencies. As throttLL’eM dynamically adjusts frequencies to meet SLOs, a wide range of power values is encountered, including peak power equal to that of Triton when under high system pressure. Overall, throttLL’eM offers 36.3% higher energy efficiency on average with oracle predictions, dropping to 30.0% at 30% prediction error level. The highest increase in energy efficiency was experienced for llama2-13b-TP2, producing up to 44.3% more tokens per Joule than Triton. Overall, energy consumption is reduced by an average of 24.7% and up to 30.7%.
(a) End-to-End latency distribution. The red dotted line denotes the E2E SLO of each model as defined in Table 2. Text boxes report the $99^{th}$ E2E latency percentile achieved by the different approaches.
@@
Figure 8: Comparison of E2E, TBT and Power consumption between Triton [triton-inference-server] and throttLL’eM. At each plot, the leftmost box represents Triton, followed by throttLL’eM with error levels of $0\%$, $15\%$, and $30\%$, respectively.
@@
These improvements are enabled by the selection of lower frequencies. The barplots of Figure 9, show that throttLL’eM selects lower, more efficient frequencies than the default 1410 MHz, ranging between 950 and 1260 MHz on average, depending on the LLM engine and the prediction error of the predictor. Higher prediction errors lead to the selection of higher frequencies, as throttLL’eM increases the predicted generation length, leading to frequency overprovisioning.
Finally, we analyze the overheads imposed by the introduction of our framework. throttLL’eM maintains a queue where queries are held until the SLOs are validated. This introduces additional latency due to the time needed to perform the validation and the possible re-queuing of a request if it is predicted to cause violations. Figure 9 contains representative examples for both situations. Llama3-8b-TP1 exhibits relatively high queuing delays as high request arrival rates stress the Scheduler, which can only schedule one request at a time. Llama2-13b-TP1 also exhibits delays, as the limited KV capacity leads to queuing until sufficient blocks become available. From the user’s perspective queuing delays add to the latency of the prefill phase (TTFT). Figure 9 shows that throttLL’eM increases TTFT when compared to Triton. A significant part of this increase is attributed to queuing, while the rest is due to the lower frequencies selected, which slow down the compute-bound prefill phase [agrawal2023sarathi].
5.4.2 throttLL’eM with autoscaling
@@
5.5 throttLL’eM Runtime Analysis
@@
In general, throttLL’eM selects lower frequencies when the active engine is under light load. When however approaching the maximum load of an active engine (max load in Table 2), frequency is increased above the energy efficiency “sweet spot” to prevent SLO violations.
@@
Figure 11 features some points of interest. The red stars mark transient E2E SLO violations due to a sudden load increase that cannot be absorbed by the TP2 engine while TP4 comes online. Yellow stars highlight how frequency is reduced when the instance is operating at lower load but cannot be downsized. Finally, green stars mark consecutive switches, as the encountered load remains close to the TP1/TP2 engine switching load boundary of 1.125 RPS.